A sample containing 223Ra (t1/2 = 11.4 days) has an activity of 8.80×105 Bq.

1st attempt Part 1 (1 point)

How many atoms of 223Ra are present?

Part 2 (1 point)Feedback How many moles of 223Ra are present?
Answer #1

We know Half life is the time taken by the nuclei to half decay.

Half time here= ln2/λ

Where λ = decay constant

Activity is the average no of disintegration per second

A= λN

Where N= the no of undecayed nuclei

So using both the equations we get

N= A× half time(in seconds)/ln2

= 8.8×10⁵ × 11.4×24×60×60/0.6931

=1.25× 10¹²

So 1.25×10¹² atoms are present.

Moles of Ra left =1.25×10¹²/6×10²³

= 2.08×10‐¹²moles